OpenStudy (anonymous):

I don't understand DeMoivre's theorem too well, nor do I feel it will be incredibly important, in PreCal or even Calculus. I have a specific question therefore that I could really use some help with: Find (1 + i)^8 using DeMoivre's Theorem. I guess it wants me to write it in standard form. This thing does a god awful job of explaining the instructions clearly. Also, if you've taken PreCal or Calculus, can you please let me know how important this theorem is? Thanks!!

OpenStudy (anonymous):

ok here is what we need to do, because demoivre is easy to implement first we have to write \(1+i\) in trigonometric form do you know how to do that?

OpenStudy (anonymous):

guess not oh well...

OpenStudy (anonymous):

the idea is to write \(i+i\) in trig form , which is straightforward enough. you need to write it as \[1+i=r(\cos(\theta)+i\sin(\theta))\] then raising it to the power of 8 will be a snap take \(r^8\) and multiply \(\theta\) by 8 that is why it is so easy, just multiply the angle

OpenStudy (anonymous):

in order to turn \(1+i\) in to \(r(\cos(\theta)+i\sin(\theta))\) you need to numbers, \(r\) and \(\theta)\) \(r=\sqrt{a^2+b^2}\) in your case \(r=\sqrt{1^2+1^2}=\sqrt{2}\) \(\theta \) is the angle where \(\tan(\theta)=\frac{b}{a}\) so in your case \(\tan(\theta)=1\) and so \(\theta=\frac{\pi}{4}\)

OpenStudy (anonymous):

*two numbers

OpenStudy (anonymous):

therefore \(1+i=\sqrt{2}(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4}))\) this in an equality that is easy to verify

OpenStudy (anonymous):

now raising it to the power of 8 is easy \(\sqrt{2}^8(\cos(\frac{\pi}{4}\times 8)+i\sin(\frac{\pi}{4}\times8))\)

OpenStudy (anonymous):

last job is to evaluate this number, and i will leave that up to you

OpenStudy (anonymous):

Thanks !! :D

OpenStudy (anonymous):

yw

OpenStudy (richyw):

you will use this theorem A LOT in calculus. To be honest I didn't even know it was a theorom, but pretty much every time you use eulers formula you will also use de Moivre's.

OpenStudy (richyw):

well I guess not pretty much everytime, but a bit still

OpenStudy (anonymous):

I forget my trig functions a lot, and Calc uses a lot of trig functions. I use it to derive the different identities/formulas from Euler's formula with de Moivre's

OpenStudy (anonymous):

can pretty much derive almost all of your trig identities by writing \(\cos(\theta)+i\sin(\theta)=e^{i\theta}\) , using the regular old laws of exponents, and then equating real and imaginary parts for example you can prove the "addition angle" formula for sine and cosine using this